OpenStudy (anonymous):

xy - 3x - 8y + 24 factor... i no you factor the x and ys, but then for 8 and 24 you find the common denominator right?

OpenStudy (anonymous):

(xy-3x)-(8y+24) idk understand what to do after besides factoring the x and ys

OpenStudy (mertsj):

xy-3x=x(y-3) -8y+24=-8(y-3)

OpenStudy (anonymous):

xy-3x-8y+24 =x(y-3)-8(y-3) =(x-8)(y-3)

OpenStudy (anonymous):

okay i get it thank you so much!!

OpenStudy (mertsj):

yw